OpenStudy (anonymous):

a particle is projected with horizontal velocity u from the top of a plane at an angle theta with the horizontal please find the distance at which the particle strikes the plane

OpenStudy (anonymous):

\[the \angle \theta \] is the angle between the plane and the horizontal

OpenStudy (anonymous):



OpenStudy (anonymous):

The x and y components are independent. Ignore x for now. \[s=ut+0.5at^2 \] \[a=-mg \] \[u=v_y(0)=v_0sin(\theta) \] \[s_y=ut+0.5at^2=v_0sin(\theta)t-0.5mgt^2 \] At hitting the ground again,s_y=0 (evidently) Rearrange to find the time Insert the time into a similar equation for s_x, which I trust you can construct by yourself (or by analogy with the s_y one).

OpenStudy (anonymous):

Note: the equation for t you will get is a quadratic, so you will obtain 2 values for t. This, again, is obvious as there are 2 moments in time when the ball touches the floor (i.e. s_y=0), when it is launched and when it lands.

OpenStudy (anonymous):

\[0 = Vy*t _{f} -\frac{ g*t _{f}^{2} }{ 2 }\] \[Vy =\frac{ g*t _{f}^{} }{ 2 }\] \[ \frac{2V _{y} }{g} =t _{f}\] \[\tan \theta =\frac{ V _{y} }{ u }\] \[distance = u*t _{f}\]

OpenStudy (anonymous):

Sorry, \[F=mg=ma\]

OpenStudy (anonymous):

How amateurish of me.

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Although not intuitively obvious, with air resistance existing and all.

OpenStudy (anonymous):

air resistance is ignored until fluid dynamics